Allison:

Jerry had $1,018.00 in savings. Last week he spent $84.00 on a desk and $372.00 on a computer from his savings. This week at a part time job sacking groceries Jerry earned $88.00 and put $44.00 of it back into savings. How much more money does Jerry need to put into savings to be back to the original balance of $1,018.00? A) $324.00 B) $368.00 C) $412.00 D) $456.00

Allison:

I got C *~* @Vocaloid if you're free, hun~

Vocaloid:

yeah that's what I got too

Allison:

On a given morning the temperature was 71°F. The temperature dropped 5°F and then rose 8°F. How much does the temperature need to rise or fall to return to the initial temperature of 71°F? A) Fall 8°F B) Fall 3°F C) Rise 8°F D) Rise 5°F

Allison:

I got B. .-.

Vocaloid:

yep, good

Allison:

What is the additive inverse of 125,000? A) −125,000 B) 125,000 C) -- 1 / 125,000 D) 1 / 125,000

Allison:

A?

Vocaloid:

yes, good job

Allison:

32 − (-19) = A) -51 B) -13 C) 13 D) 51

Allison:

D?

Vocaloid:

yes

Allison:

Oh you starred me :) Thank chu c:

Allison:

National Oil Company reported a net loss last year of $14 million. This year their net gain is $65 million. How much more money did they make this year than last year? A) $79 million B) $65 million C) $14 million D) $51 million

Allison:

D?

Vocaloid:

a net loss means a negative number, so (-14 million) 65 million - (-14 million) = ?

Allison:

Oh o: $79 Million?

Vocaloid:

yup

Allison:

6 + (-27) = A) -33 B) -21 C) 21 D) 33

Allison:

-21?

Vocaloid:

yup, good

Allison:

22 − (-18) = A) -40 B) -4 C) 4 D) 40

Allison:

D

Vocaloid:

right

Allison:

What is the sum of 14 and the opposite of 12? A) -26 B) -2 C) 2 D) 26

Allison:

26? >.<

Allison:

Gahh.

Vocaloid:

whoops

Vocaloid:

sorry it should be 14 + (-12)

Allison:

2?

Vocaloid:

good, so C

Allison:

Simplify |-5| + |6|. A) -30 B) -1 C) 1 D) 11

Allison:

C?

Vocaloid:

these things | | are absolute value symbols you just "erase" the negative signs inside the symbols and then treat them like regular numbers

Vocaloid:

so |-5| becomes 5 and |6| is just 6 since it wasn't negative

Allison:

11? ._.

Vocaloid:

so |-5| + |6| = 5 + 6 = 11 good

Allison:

-14.25 + -3.78 - (-5.78) - 3.4 A) -19.65 B) -15.65 C) 1.29 D) 27.21

Allison:

B?

Vocaloid:

yup good

Allison:

A mountain climber starts a climb at an elevation of 453 feet above sea level. At his first rest stop he has climbed 162 feet, and by his second rest stop he has climbed another 207 feet. It's getting late in the day, so the climber starts his way down. If the climber descends 285 feet, how much does he need to ascend or descend to return to the original starting point? A) ascend 84 feet B) ascend 78 feet C) descend 84 feet D) descend 78 feet

Allison:

I wanna say C..?

Vocaloid:

yup good

Allison:

A submarine is 671 feet below sea level. If it ascends 218 feet and then descends 1363 feet, what is its new position? A) 474 feet below sea level B) 1,363 feet below sea level C) 1,816 feet below sea level D) 2,252 feet below sea level

Allison:

I dunno >.<

Vocaloid:

"ascends" means it's going up in height "descends" means it's going down in height

Vocaloid:

so "If it ascends 218 feet and then descends 1363 feet" means you are adding 218 feet and subtracting 1363 from the original height 671

Vocaloid:

so the calculation is: 671 + 218 - 1363 = ?

Allison:

A..?

Vocaloid:

yup A

Vocaloid:

(I should have wrote a negative sign in front of the 671 whoops)

Allison:

Same answer anyway xD

Allison:

The elevation of the basement floor in a building is -14 ft. The elevation of the roof is 36 feet. What is the distance from the basement floor to the roof? A) -50 ft B) -24 ft C) 24 ft D) 50 ft

Allison:

A?

Vocaloid:

you're on the right track, but we usually write distances as positive numbers only, so 50

Allison:

21 + (-33) = A) -54 B) -12 C) 12 D) 54

Allison:

I got B *~*

Vocaloid:

good

Allison:

-18 − (-5) = A) -23 B) -13 C) 13 D) 23

Allison:

B?

Vocaloid:

good

Allison:

So, I checked on 15 questions with you and got one wrong *~*

Vocaloid:

oh dear do you know which one ;;

Allison:

It didn't show :c

Vocaloid:

oh whoops I think it was the submarine one ;;

Allison:

It's okaii doe, I'll live c: Thank you for the halp <3

Allison:

Oh o:

Vocaloid:

should have been -671 + 218 - 1363 = - 1816 not positive 671

Allison:

._. Oh.

Allison:

That's enough math for the night lol

Allison:

Thank you ^-^